OpenStudy (anonymous):

Indicate which of the following soil pH ranges: A. acidic (pH 4 to 6.5), B. basic (7.2 to 9.0), and C. near neutral (pH 6.5 to 7.2) best apply to the scenario below: hydroxyl ion (OH-) concentration equal to 3 x 10-6 molar a. acidic b. basic c. neutral

OpenStudy (caozeyuan):

POH=-log(3*10^-6)

OpenStudy (caozeyuan):

pOH=5.5,Ph=14-5.5=8.5

OpenStudy (anonymous):

thanks